OpenStudy (anonymous):

Temperatures in degrees Fahrenheit can be converted to temperatures in degrees Celsius using the formula c=(5/9) F-32 degrees What is the temperature in degrees Celsius if an outdoor thermometer reads 59°F?

OpenStudy (anonymous):

you got it? pleassse

OpenStudy (anonymous):

you have the formula incorrect.. just a minute

OpenStudy (anonymous):

c=(5/9)(f-32) c=(5/9)(59-32)=15 degrees c

OpenStudy (anonymous):

thnks

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

try A function is shown in the table below. Determine the rule for the function. Input, x Output, y 2 13 5 31 7 43 9 55